find the x and y intercepts and graph 2x+3y=4
Answer by checkley75(3666)

2x+3y=4
3y=-2x+4
y=-2x/3+4/3 set y=0 and solve for x.
0=-2x/3+4/3
-2x/3=-4/3 cross multiply.
-6x=-12
x=-12/-6
x=2 solution this point is (2,0).
setting x=0 then we solve for y.
y=-2*0/3+4/3
y=4/3 solution (0,4/3).